Molly is selling bracelets and necklaces at a craft fair. The cost of 4 bracelets and 3 necklaces is $23.50. The cost of 5 bracelets and 2 necklaces is $21.50. The system of equations shown below represents this situation.
4b + 3n = $23.50
5b + 2n = $21.50
A customer bought 1 bracelet and 1 necklace. How much money did the customer spend?

Answer by FrankM(1040)

4b + 3n = $23.50
5b + 2n = $21.50
let's multiply the first equation by 2, and the second by 3:
8b + 6n = $47.00
15b + 6n = $64.50
subtract 1 from 2:
7b = $17.50
b = $2.50
back to first equation:
4($2.50) +3n = $23.50
3n = $13.50
n= $4.50
one of each total $7.00